The glass core of an optical fiber has an index of refraction 1.60. The index of refraction of the cladding is 1.49.


A) What is the maximum angle a light ray can make with the wall of the core if it is to remain inside the fiber?
Express your answer with the appropriate units.
Answer #1

From Snell's Law we have n1 * sin(theta) = n2 * sin(90) = n2

so theta = arcsin(n2/n1)


= arcsin(1.49 / 1.60) = 68.63 degree


But this angle is measured from the normal so from the wall the angle


= 90 -68.63 = 21.37 degree